# Masked Language Modeling Becomes Conditional Density Estimation for Tabular Data Synthesis

### v1

@@ Abstract
In this paper, our goal is to generate synthetic data for heterogeneous (mixed-type) tabular datasets with high machine learning utility (MLu). Given that the MLu performance relies on accurately approximating the conditional distributions, we focus on devising a synthetic data generation method based on conditional distribution estimation. We propose a novel synthetic data generation method, MaCoDE, by redefining the multi-class classification task of Masked Language Modeling (MLM) as histogram-based non-parametric conditional density estimation. Our proposed method enables estimating conditional densities across arbitrary combinations of target and conditional variables. Furthermore, we demonstrate that our proposed method bridges the theoretical gap between distributional learning and MLM. To validate the effectiveness of our proposed model, we conduct synthetic data generation experiments on 10 real-world datasets. Given the analogy between predicting masked input tokens in MLM and missing data imputation, we also evaluate the performance of multiple imputations on incomplete datasets with various missing data mechanisms. Moreover, our proposed model offers the advantage of enabling adjustments to data privacy levels without requiring re-training.
@@ 1 Introduction
There are two main objectives in synthetic data generation: (1) preserving the statistical characteristics of the original dataset and (2) achieving comparable machine learning utility (MLu) to the original dataset. However, achieving high statistical fidelity does not guarantee high MLu performance [15], which implies that the approach to synthetic data generation should align with the intended goal. In this paper, our focus is on generating synthetic data with high MLu performance. Given that MLu performance relies on accurately approximating the conditional distribution, we concentrate on devising a synthetic data generation method based on estimating the conditional distribution.
Thus, our primary contribution lies in introducing a novel masked learning method for the conditional density estimation of heterogeneous (mixed-type) tabular datasets. Our proposed method leverages the approach of [29] and Masked Language Modeling (MLM) [8] and diverges from existing methodologies, which focus on minimizing the discrepancy between the ground-truth distribution and the generative model [51, 56, 1, 24]. Note that [29] demonstrated that the conditional density estimation problem can be transformed into a multi-class classification problem.
MLM involves randomly masking a portion of input tokens during training to predict the original words based on their context. In the natural language domain, previous attempts to interpret MLM as distributional learning have been somewhat limited, relying on pseudo-likelihood or Markov random fields [12, 49, 42, 38, 17]. However, we redefine the multi-class classification task of MLM as histogram-based non-parametric conditional density estimation and show that our proposed method bridges the theoretical gap between distributional learning and the MLM approach. We term our proposed model MaCoDE (Masked Conditional Density Estimation).
@@
By adopting a non-parametric approach for conditional density estimation, our proposed model addresses the challenge of modeling non-uniform distributions of continuous columns [28, 11]. Since the histogram-based approach is theoretically valid only when the continuous variables have bounded supports [50, 29], we transform continuous columns using the Cumulative Distribution Function (CDF) and constrain their values to the interval $[0,1]$ [28, 11].
Furthermore, by employing a masking scheme where the number of masked tokens is uniformly distributed and the model structure of BERT (masked transformer encoder) [8], our proposed model enables the estimation of conditional densities across arbitrary combinations of target and conditional variables [12, 19, 37]. It leads to our proposed model to generate columns of tabular data in a random order, reflecting the inherent property of tabular data not having intrinsic ordering [13]. This contrasts with the existing auto-regressive density estimators, where the generation procedure proceeds in a fixed order of columns [14, 22, 27].
Meanwhile, incomplete tabular data poses a widespread challenge in real-world data collection [47]. In MLM, the masked input tokens can be seen as missing data, and predicting the original words can be analogized to imputation. In this context, [13] proposed TabMT, an MLM-based synthesizer capable of handling incomplete data. However, since TabMT is based on predicting the K-means cluster index of masked entries, it encounters challenges for distributional learning. Conversely, our proposed model can estimate the conditional distribution while simultaneously accommodating arbitrary conditioning sets to address diverse missingness patterns - an essential capability for generating samples and performing multiple imputations [47, 19, 37]. We substantiate the effectiveness of our proposed method by evaluating its performance in both synthetic data generation and multiple imputations. This evaluation encompasses 10 real-world tabular datasets and various missing data mechanisms.
@@ 2 Proposal
Figure 1: Overall structure and training process of MaCoDE. In this case, the value of the second column is masked (replaced with ‘0’) and predicted.
Notations. Let ${\bf x}\in{\cal X}_{1}\times\cdots\times{\cal X}_{p}\subseteq\mathbb{R}^{p}$ denote an observation consisting of continuous and categorical (discrete) variables, where ${\cal X}_{j}$ denotes the support of $j$th variable (column), ${\bf x}_{j}$. Here, subscript $j$ refers to the $j$th element. $I_{C}$ and $I_{D}$ represent the index sets for continuous and categorical variables, where $I_{C}\cup I_{D}=\{1,\cdots,p\}$. The observed dataset is denoted as $\{{\bf x}^{(i)}\}_{i=1}^{n}$. ${\bf m}\in\{0,1\}^{p}$ is a binary vector indicating masked values, with ${\bf m}_{j}=0$ indicating the $j$th column is masked (if ${\bf m}_{j}=1$, then the $j$th column is not masked). $F^{*}_{j}$ indicates the ground-truth CDF of the $j$th column, and we assume Assumption 1.
@@
For all $j\in I_{C}$, $F^{*}_{j}:\mathbb{R}\mapsto[0,1]$ is continuous, differentiable, and strictly monotonically increasing.
2.1 Classification Target (Discretization)
@@
Continuous column. Firstly, we transform the continuous variables into random variables within the $[0,1]$ range using their marginal CDFs [28, 11]. Then, we partition the $[0,1]$ interval with $L+1$ cut-points, $b_{0},b_{1},\cdots,b_{L}$, where $10=b_{0}<b_{1}<b_{2}<\cdots<b_{L-1}<b_{L}=$ 1, resulting in $L$ bins. We define the classification target ${\bf y}_{j}$ based on the interval within which $F^{*}_{j}({\bf x}_{j})$ falls, as follows:
${\bf y}_{j}\coloneqq\sum_{s=0}^{L-1}\mathbb{I}\Big{(}b_{s}\leq F^{*}_{j}({\bf x}_{j})\Big{)},$
indicating that the classification target is the bin index, where $\mathbb{I}(\cdot)$ represents the indicator function, and ${\bf y}_{j}\in[L]\coloneqq\{1,2,\cdots,L\}$. Thus, if $b_{l-1}\leq F^{*}_{j}({\bf x}_{j})<b_{l}$, then ${\bf y}_{j}=l$.
Our discretization procedure presents an advantage over [29], where bins are structured along the real line. Since the lower and upper bounds are naturally defined as 0 and 1, respectively, within our approach, determining the number of bins becomes more intuitive. For instance, one can use finer bins at the boundaries of 0 and 1 to achieve more accurate tail density estimation.
@@
The discretization function $g:{\cal X}_{1}\times\cdots\times{\cal X}_{p}\mapsto[L]^{p}$ is defined as
$g({\bf x};F^{*})_{j}$$\coloneqq$ $\begin{cases}\sum_{s=0}^{L-1}\mathbb{I}(b_{s}\leq F^{*}_{j}({\bf x}_{j})),&\text{if $j\in I_{C}$}\\ {\bf x}_{j},&\text{if $j\in I_{D}$}\end{cases},$
for $j=1,\cdots,p$, where $F^{*}\coloneqq\{F^{*}_{j}:j\in I_{C}\}$.
Finally, we transform the observations ${\bf x}$ into the discretized label ${\bf y}$ using the discretization function $g$ of Definition 1, as ${\bf y}_{j}=g({\bf x};F^{*})_{j}$ for all $j$.
2.2 Masked Conditional Density Estimation (MaCoDE)
Ground-truth distribution. For all $j\in\{1,2,\cdots,p\}$, the ground-truth conditional probability of $g({\bf x};F^{*})_{j}=l$ given the other observed variables, i.e., $\{{\bf x}_{k}:{\bf m}_{k}=1,k\neq j\}$, is defined as
$\Pr\left(g({\bf x};F^{*})_{j}=l\Big{|}\{{\bf x}_{k}:{\bf m}_{k}=1,k\neq j\}\right)\eqqcolon\pi^{*}_{jl}\Big{(}\{{\bf x}_{k}:{\bf m}_{k}=1,k\neq j\}\Big{)},$
which also corresponds to the conditional probability of $F_{j}^{*}({\bf x}_{j})$ in the $l$th bin. Then, the ground-truth conditional distribution of $g({\bf x};F^{*})_{j}$ is written accordingly:
$p_{j}^{*}\Big{(}g({\bf x};F^{*})_{j}|\{{\bf x}_{k}:{\bf m}_{k}=1,k\neq j\}\Big{)}=\prod_{l=1}^{L}\pi^{*}_{jl}\Big{(}\{{\bf x}_{k}:{\bf m}_{k}=1,k\neq j\}\Big{)}^{\mathbb{I}(g({\bf x};F^{*})_{j}=l)}.$ (1)
@@ 2.2 Masked Conditional Density Estimation (MaCoDE)
$p_{j}\left(g({\bf x};\hat{F})_{j}|g({\bf x};\hat{F})\odot{\bf m};\theta\right)=\prod_{l=1}^{L}\pi_{jl}\left(g({\bf x};\hat{F})\odot{\bf m};\theta\right)^{\mathbb{I}(g({\bf x};\hat{F})_{j}=l)},$
where $\odot$ denotes element-wise multiplication, $\sum_{l=1}^{L}\pi_{jl}(g({\bf x};\hat{F})\odot{\bf m};\theta)=1$ for all $j$, $\hat{F}\coloneqq\{\hat{F}_{j}:j\in I_{C}\}$, and $\hat{F}_{j}$ is an estimator of $F^{*}_{j}$. In this paper, we use the empirical CDF for $\hat{F}_{j}$. Here, $\theta$ represents all the parameters of the transformer encoder-based classifier [8], allowing us to process inputs of arbitrary lengths and accommodate different combinations of observed and masked variables. Additionally, we denote $(\pi_{j1}(\cdot;\theta),\cdots,\pi_{jL}(\cdot;\theta))$ as $\pi_{j}(\cdot;\theta)$.
Then, the objective function for a single observation is defined as the log-likelihood of masked entries:
${\cal L}({\bf y},{\bf m};\theta)\coloneqq\sum_{j:{\bf m}_{j}=0}\log p_{j}({\bf y}_{j}|{\bf y}\odot{\bf m};\theta)=\sum_{j:{\bf m}_{j}=0}\sum_{l=1}^{L}\mathbb{I}({\bf y}_{j}=l)\cdot\log\pi_{jl}({\bf y}\odot{\bf m};\theta),$ (2)
where the label ${\bf y}_{j}$ is defined as ${\bf y}_{j}=g({\bf x};\hat{F})_{j}$ for all $j$. Our objective function 2 estimates the conditional distribution of $\hat{F}_{j}({\bf x}_{j})$ in each bin, where its target distribution is (1). Similar to MLM, $({\bf y}\odot{\bf m})_{j}$ serves as a ‘word’ (bin) index aligned with the $j$th column, encompassing its own vocabulary set of $L+1$ words (=$=\{0,1,2,\cdots,L\}$), including ‘0’ for the masked input. And $\pi_{jl}$ represents the probability that the model outputs the word index $l$ from the $j$th column.
For each $j$, the embedding layer preprocesses of $({\bf y}\odot{\bf m})_{j}$ through one-hot encoding, with each bin index being assigned a learnable embedding vector. Note that our approach to handling masked entries shares similarities with the existing zero imputation technique, where all masked entries are replaced with zeros [19, 34, 37, 18]. However, in our proposed method, although all masked entries are replaced with the same bin index, ‘0’, each imputed zero value is embedded through distinct embedding vectors for each column.
Definition 2 (Mask distribution [13, 19]).
@@
where the discretized training dataset is $\{{\bf y}^{(i)}\}_{i=1}^{n}=\{g({\bf x}^{(i)};\hat{F})\}_{i=1}^{n}$, and $\mathbb{E}_{p({\bf m})}$ is approximated by Monte-Carlo and ancestral sampling. The distribution $p({\bf m})$ can be defined by the user based on the specific problem requirements [19]. As in BERT [8] and TabMT [13], the task of our objective function (3) is to predict the original label for all masked inputs. The overall structure and the training procedure of MaCoDE are outlined in Figure 1 and Algorithm 1, respectively (note that mini-batch style training is also applicable).
@@
We want to emphasize that our objective function (3) does not imply an assumption of conditional independence among masked entries given the observed data. Instead, by ensuring that $p({\bf m})$ has full support over $\{0,1\}^{p}$, sampling ${\bf m}$ from $p({\bf m})$ allows us to learn conditional densities encompassing all possible combinations of conditioning sets and target variables [19, 13]. Furthermore, this training scheme remains invariant to the missing data scenario.
Synthetic data generation. Tabular data lacks the inherent ordering between columns, unlike natural language [13]. Therefore, as outlined in Algorithm 2, MaCoDE randomly generates one column at a time, conditioned on masked subset sizes from $p$ to $1$, in descending order ($p\rightarrow p-1\rightarrow\cdots\rightarrow 2\rightarrow 1$). [13] demonstrated that, under the masked distribution of Definition 2, the distribution of the number of masked entries is matched during both training and generation.
Figure 2: Trade-off between quality and privacy. Left: feature selection performance. Right: DCR. Error bars represent standard errors. See Appendix A.6.1 for detailed results.
Controllable privacy level. Adjusting the privacy level during synthetic data generation is crucial in tabular domain [39]. Similar to TabMT [13], we can also regulate the privacy level using a single hyper-parameter $\tau$ without the need for re-training, while other existing synthesizers have fixed trade-offs between synthetic data quality and privacy level [57, 24] or require re-training [1]. By increasing the temperature parameter $\tau$ in $\hat{{\bf y}}_{j}\sim Cat\big{(}\pi_{j}(\hat{{\bf y}}\odot\hat{{\bf m}};\theta)/\tau\big{)}$ at line 2 of Algorithm 2, we can mitigate the risk of privacy leakage. Figure 2 shows that MaCoDE allows for a trade-off between synthetic data quality (feature selection performance) and privacy preservability (DCR, Distance to Closest Record) as $\tau$ increases.
Algorithm 1 MaCoDE
Input: $\mathcal{D}\coloneqq\{{\bf x}^{(i)}\}_{i=1}^{n}$; $g$.
1:For $j\in I_{C}$, estimate $\hat{F}_{j}$ using $\mathcal{D}$
2:$\{{\bf y}^{(i)}\}_{i=1}^{n}=\{g({\bf x}^{(i)};\hat{F})\}_{i=1}^{n}$
3:for $e=1,2,\cdots,epochs$ do
4:     ${\bf m}^{(i)}\sim_{i.i.d.}p({\bf m})$ for $i=1,\cdots,n$
5:     Update $\theta$ by the gradient descent of $\sum_{i=1}^{n}{\cal L}({\bf y}^{(i)},{\bf m}^{(i)};\theta)$
@@
Initialize: For all $j$, $\hat{{\bf y}}_{j}\leftarrow 0$ and $\hat{{\bf m}}_{j}\leftarrow 0$.

### v2

@@ Abstract
In this paper, our goal is to generate synthetic data for heterogeneous (mixed-type) tabular datasets with high machine learning utility (MLu). Since the MLu performance depends on accurately approximating the conditional distributions, we focus on devising a synthetic data generation method based on conditional distribution estimation. We introduce MaCoDE by redefining the consecutive multi-class classification task of Masked Language Modeling (MLM) as histogram-based non-parametric conditional density estimation. Our approach enables the estimation of conditional densities across arbitrary combinations of target and conditional variables. We bridge the theoretical gap between distributional learning and MLM by demonstrating that minimizing the orderless multi-class classification loss leads to minimizing the total variation distance between conditional distributions. To validate our proposed model, we evaluate its performance in synthetic data generation across 10 real-world datasets, demonstrating its ability to adjust data privacy levels easily without re-training. Additionally, since masked input tokens in MLM are analogous to missing data, we further assess its effectiveness in handling training datasets with missing values, including multiple imputations of the missing entries.
@@ 1 Introduction
There are two main objectives in synthetic tabular data generation: (1) preserving the statistical characteristics of the original dataset and (2) achieving comparable machine learning utility (MLu) to the original dataset. In this paper, our focus is on generating synthetic data with high MLu performance. Note that achieving high statistical fidelity does not guarantee high MLu performance (Hansen et al. 2023).
Given that MLu performance depends on accurately approximating conditional distributions, we focus on developing a synthetic data generation method based on conditional distribution estimation. However, two important properties of tabular data must be considered: ($i$) tabular data can consist of mixed types of data (Borisov et al. 2021; Shwartz-Ziv and Armon 2022), and ($ii$) the tabular data does not have an intrinsic ordering among columns (Gulati and Roysdon 2023).
@@
($i$) Considering the heterogeneous nature of tabular data and aiming to develop a method that addresses the challenges of modeling diverse distributions of continuous columns, we employ histogram-based non-parametric conditional density estimation through a multi-class classification task (Li, Bondell, and Reich 2019). This approach enables us to apply the classification loss uniformly across all types of columns. Since the histogram-based approach is theoretically valid only when continuous variables have bounded supports (Wasserman 2006; Li, Bondell, and Reich 2019), we transform continuous columns using the Cumulative Distribution Function (CDF) and constrain their values to the interval $[0,1]$ (Li et al. 2021; Fang et al. 2022).
($ii$) To learn the arbitrary generation ordering of columns, we utilize the Masked Language Modeling (MLM) approach (Devlin et al. 2019). By employing a masking scheme and the BERT model architecture, our proposed model enables the estimation of conditional densities across arbitrary combinations of target and conditional variables (Ghazvininejad et al. 2019; Ivanov, Figurnov, and Vetrov 2019; Nazábal et al. 2020). (Gulati and Roysdon 2023) proposed a similar method called TabMT, however, TabMT faces challenges in distributional learning because it relies on predicting the K-means cluster index of masked entries. Our approach contrasts with existing auto-regressive density estimators, which generate data in a fixed column order (Hansen 1994; Kamthe, Assefa, and Deisenroth 2021; Letizia and Tonello 2022). Additionally, (Germain et al. 2015; Papamakarios, Pavlakou, and Murray 2017) are also able to estimate conditional densities but differ from our approach by masking the model weights rather than the input.
Therefore, our proposed method redefines the consecutive multi-class classification task of MLM as histogram-based non-parametric conditional density estimation. We term our proposed model MaCoDE (Masked Conditional Density Estimation). The main contribution of our work is bridging the theoretical gap between distributional learning and the consecutive minimization of the multi-class classification loss within the MLM approach.
Specifically, we demonstrate that minimizing the orderless multi-class classification loss, when combined with the CDF transformation, provides theoretical validity for minimizing the discrepancy between conditional distributions in terms of total variation distance. This implies that we do not need to consider the ordering of bins, which could otherwise serve as a useful inductive bias. Note that, in the natural language domain, previous attempts to interpret MLM as distributional learning have been somewhat limited, relying on pseudo-likelihood or Markov random fields (Ghazvininejad et al. 2019; Wang and Cho 2019; Salazar et al. 2019; Ng, Cho, and Ghassemi 2020; Hennigen and Kim 2023).
We substantiate the effectiveness of our proposed method by evaluating its performance in synthetic data generation across 10 real-world tabular datasets, demonstrating its capability to adjust data privacy levels easily without re-training. Given that masked input tokens in MLM can be viewed as missing data, we also assess our model’s effectiveness in handling training datasets with missing values. Moreover, as our proposed model estimates the conditional distribution while accommodating arbitrary conditioning sets, it can address various missingness patterns - an essential capability for generating samples and performing multiple imputations (van Buuren 2012; Ivanov, Figurnov, and Vetrov 2019; Nazábal et al. 2020). Consequently, we further validate our method’s effectiveness by evaluating its performance in multiple imputations across various missing data mechanisms.
2 Related Works
Existing methods using deep generative models aim to directly minimize the discrepancy between the multivariate ground-truth distribution and the generative model. These include CTGAN (Xu et al. 2019), TVAE (Xu et al. 2019), CTAB-GAN (Zhao et al. 2021), CTAB-GAN+ (Zhao et al. 2023b), DistVAE (An and Jeon 2023), and TabDDPM (Kotelnikov et al. 2023).
In the realm of transformer-based synthesizers, methods such as TabPFGen (Ma et al. 2023), TabMT (Gulati and Roysdon 2023), and REaLTabFormer (Solatorio and Dupriez 2023) have been proposed. TabMT utilizes an MLM-based approach to generate synthetic data by predicting cluster indices from K-means clustering. TabPFGen is an energy-based model that leverages the Bayesian inference of TabPFN (Müller et al. 2022) framework. REaLTabFormer employs an autoregressive Transformer architecture akin to GPT-2 (Radford et al. 2019), applying natural language generation techniques to tabular data. Additionally, methods such as (Kamthe, Assefa, and Deisenroth 2021; Letizia and Tonello 2022; Drouin, Marcotte, and Chapados 2022; Ashok et al. 2024) use transformers for copula density estimation, specifically in time-series datasets.
@@ 3 Proposal
Figure 1: Overall structure of MaCoDE. In this case, the value of the second column is masked (replaced with ‘0’) and predicted.
Notations. Let ${\bf x}\in\mathbb{R}^{p}$ denote an observation consisting of continuous and categorical (discrete) variables, and the $j$th variable (column) is denoted as ${\bf x}_{j}$. Here, subscript $j$ refers to the $j$th element. $I_{C}$ and $I_{D}$ represent the index sets for continuous and categorical variables, where $I_{C}\cup I_{D}=\{1,\cdots,p\}$. The observed dataset is denoted as $\{{\bf x}^{(i)}\}_{i=1}^{n}$. ${\bf m}\in\{0,1\}^{p}$ is a binary vector indicating masked values, with ${\bf m}_{j}=0$ indicating the $j$th column is masked (if ${\bf m}_{j}=1$, then the $j$th column is not masked). $F^{*}_{j}$ indicates the ground-truth CDF of the $j$th column, and $\hat{F}_{j}$ is an estimator of $F^{*}_{j}$.
Overview. Without loss of generality, we can consider an arbitrary conditional density function: for $j\in I_{C}$,
$p_{j}^{*}({\bf x}_{j}|{\bf x}_{1},\cdots,{\bf x}_{j-1}).$ (1)
Our primary objective is to estimate (1). However, there are two major challenges in estimating (1):
  1. 1.
    Modeling non-uniform distributions of continuous columns, ${\bf x}_{j}$.
  2. 2.
    Handling arbitrary combinations of conditional variables, ${\bf x}_{1},\cdots,{\bf x}_{j-1}$.
@@
For all $j\in I_{C}$, there exists $\hat{F}_{j}:\mathbb{R}\mapsto[0,1]$ such that $\hat{F}_{j}$ is invertible and differentiable.
In this paper, we address these major challenges by unifying a histogram-based conditional density estimation and the MLM approach. And we assume that $\hat{F}_{j}$ satisfying Assumption 1 is given. For $j\in I_{C}$,
$p_{j}^{*}({\bf x}_{j}|{\bf x}_{-j})$ $=$ $c^{*}_{j}(\hat{F}_{j}({\bf x}_{j})|{\bf x}_{-j})\cdot\hat{p}_{j}({\bf x}_{j})$ (2)
by the change of variable in terms of $\hat{F}_{j}$, where $c^{*}_{j}$ is the conditional density of $\hat{F}_{j}({\bf x}_{j})$, ${\bf x}_{-j}\coloneqq({\bf x}_{1},\cdots,{\bf x}_{j-1})$, and $\hat{p}_{j}$ denotes the density of $\hat{F}_{j}$. In particular, we estimate $c^{*}_{j}$ in (2) using a histogram-based approach.
Classification Target (Discretization)
@@
Continuous column. Firstly, since the histogram-based approach requires continuous columns to have bounded supports, we transform them into random variables within the $[0,1]$ range using their marginal CDFs. Then, we partition the $[0,1]$ interval with $L+1$ cut-points, $b_{0},b_{1},\cdots,b_{L}$, where $10=b_{0}<b_{1}<b_{2}<\cdots<b_{L-1}<b_{L}=$ 1, resulting in $L$ bins. We define the classification target ${\bf y}_{j}$ based on the interval within which $\hat{F}_{j}({\bf x}_{j})$ falls, as follows:
${\bf y}_{j}\coloneqq\sum_{s=0}^{L-1}\mathbb{I}\Big{(}b_{s}\leq\hat{F}_{j}({\bf x}_{j})\Big{)},$
indicating that the classification target is the bin index, where $\mathbb{I}(\cdot)$ represents the indicator function, and ${\bf y}_{j}\in[L]\coloneqq\{1,2,\cdots,L\}$. Thus, if $b_{l-1}\leq\hat{F}_{j}({\bf x}_{j})<b_{l}$, then ${\bf y}_{j}=l$.
Our discretization procedure presents an advantage: since the lower and upper bounds are naturally defined as 0 and 1 within our approach, determining the number of bins becomes more intuitive. For example, finer bins near the boundaries of 0 and 1 can be employed to achieve more accurate tail density estimation.
@@
The discretization function $g:\mathbb{R}^{p}\mapsto[L]^{p}$ is defined as
$g({\bf x};\hat{F})_{j}$$\coloneqq$ $\begin{cases}\sum_{s=0}^{L-1}\mathbb{I}(b_{s}\leq\hat{F}_{j}({\bf x}_{j})),&\text{if $j\in I_{C}$}\\ {\bf x}_{j},&\text{if $j\in I_{D}$}\end{cases},$
for $j=1,\cdots,p$, where $\hat{F}\coloneqq\{\hat{F}_{j}:j\in I_{C}\}$.
Finally, we transform the observations ${\bf x}$ into the discretized label ${\bf y}$ using the discretization function $g$ of Definition 1, as ${\bf y}_{j}=g({\bf x};\hat{F})_{j}$ for all $j$.
Masked Conditional Density Estimation
Target distribution. For all $j\in\{1,2,\cdots,p\}$, the ground-truth conditional probability of $g({\bf x};\hat{F})_{j}=l$ given the other observed variables, i.e., $\{{\bf x}_{k}:{\bf m}_{k}=1,k\neq j\}$, is defined as
$\Pr\left(g({\bf x};\hat{F})_{j}=l\Big{|}\{{\bf x}_{k}:{\bf m}_{k}=1,k\neq j\}\right)$
$\eqqcolon$ $\pi^{*}_{jl}\Big{(}\{{\bf x}_{k}:{\bf m}_{k}=1,k\neq j\}\Big{)},$
which also corresponds to the conditional probability of $\hat{F}_{j}({\bf x}_{j})$ in the $l$th bin, $[b_{l-1},b_{l})$.
Then, the ground-truth conditional distribution of $g({\bf x};\hat{F})_{j}$, which is our target distribution, is written accordingly:
$p_{j}^{*}\Big{(}g({\bf x};\hat{F})_{j}|\{{\bf x}_{k}:{\bf m}_{k}=1,k\neq j\}\Big{)}$ (3)
$=$ $\prod_{l=1}^{L}\pi^{*}_{jl}\Big{(}\{{\bf x}_{k}:{\bf m}_{k}=1,k\neq j\}\Big{)}^{\mathbb{I}(g({\bf x};\hat{F})_{j}=l)}.$
@@
$p_{j}\left(g({\bf x};\hat{F})_{j}|g({\bf x};\hat{F})\odot{\bf m};\theta\right)$
$=$ $\prod_{l=1}^{L}\pi_{jl}\left(g({\bf x};\hat{F})\odot{\bf m};\theta\right)^{\mathbb{I}(g({\bf x};\hat{F})_{j}=l)},$
where $\odot$ denotes element-wise multiplication and $\sum_{l=1}^{L}\pi_{jl}(g({\bf x};\hat{F})\odot{\bf m};\theta)=1$ for all $j$. In this paper, we use the empirical CDF for $\hat{F}_{j}$. Here, $\theta$ represents all the parameters of the transformer encoder-based classifier, allowing us to process inputs of arbitrary lengths and accommodate different combinations of observed and masked variables. Additionally, we denote $(\pi_{j1}(\cdot;\theta),\cdots,\pi_{jL}(\cdot;\theta))$ as $\pi_{j}(\cdot;\theta)$.
Then, the objective function for a single observation is defined as the negative log-likelihood of masked entries:
${\cal L}({\bf y},{\bf m};\theta)$$\coloneqq$ $-\sum_{j:{\bf m}_{j}=0}\log p_{j}({\bf y}_{j}|{\bf y}\odot{\bf m};\theta)$
$=$ $-\sum_{j:{\bf m}_{j}=0}\sum_{l=1}^{L}\mathbb{I}({\bf y}_{j}=l)\cdot\log\pi_{jl}({\bf y}\odot{\bf m};\theta),$
where the label ${\bf y}_{j}$ is defined as ${\bf y}_{j}=g({\bf x};\hat{F})_{j}$ for all $j$. Our objective function (3) estimates the conditional distribution of $\hat{F}_{j}({\bf x}_{j})$ in each bin, where its target distribution is (3). In other words, our objective is to approximate the true conditional probability $\pi^{*}_{jl}$ by our estimated probability $\pi_{jl}$.
Similar to MLM, the bin index $({\bf y}\odot{\bf m})_{j}$ serves as a ‘word’ index aligned with the $j$th column, encompassing its own vocabulary set of $L+1$ words (i.e., $\{0,1,2,\cdots,L\}$), including ‘0’ for the masked input. And $\pi_{jl}$ represents the probability that the model outputs the word index $l$ from the $j$th column.
For each $j$, the embedding layer preprocesses of $({\bf y}\odot{\bf m})_{j}$ through one-hot encoding, with each bin index being assigned a learnable embedding vector. Note that our approach to handling masked entries shares similarities with the existing zero imputation technique, where all masked entries are replaced with zeros (Ivanov, Figurnov, and Vetrov 2019; Mattei and Frellsen 2019; Nazábal et al. 2020; Ipsen, Mattei, and Frellsen 2021). However, in our proposed method, although all masked entries are replaced with the same bin index, ‘0’, each imputed zero value is embedded through distinct embedding vectors for each column.
Definition 2 (Mask distribution).
@@
where the discretized training dataset is $\{{\bf y}^{(i)}\}_{i=1}^{n}=\{g({\bf x}^{(i)};\hat{F})\}_{i=1}^{n}$, and $\mathbb{E}_{p({\bf m})}$ is approximated by Monte-Carlo and ancestral sampling. The distribution $p({\bf m})$ can be defined by the user based on the specific problem requirements (Ivanov, Figurnov, and Vetrov 2019). As in BERT (Devlin et al. 2019), the task of our objective function (5) is to predict the original label for all masked inputs. The overall structure of MaCoDE is outlined in Figure 1.
@@
We want to emphasize that our objective function (5) does not imply an assumption of conditional independence among masked entries given the observed data. Instead, by ensuring that $p({\bf m})$ has full support over $\{0,1\}^{p}$, sampling ${\bf m}$ from $p({\bf m})$ allows us to learn conditional densities encompassing all possible combinations of conditioning sets and target variables (Ivanov, Figurnov, and Vetrov 2019; Gulati and Roysdon 2023). Furthermore, this training scheme remains invariant to the missing data scenario.
@@
Initialize: For all $j$, $\hat{{\bf y}}_{j}\leftarrow 0$ and $\hat{{\bf m}}_{j}\leftarrow 0$. (all columns are masked)